# EmoSphere-SER: Enhancing Speech Emotion Recognition Through Spherical Representation with Auxiliary Classification

### v1

@@ Abstract
Speech emotion recognition predicts a speaker’s emotional state from speech signals using discrete labels or continuous dimensions such as arousal, valence, and dominance (VAD). We propose EmoSphere-SER, a joint model that integrates spherical VAD region classification to guide VAD regression for improved emotion prediction. In our framework, VAD values are transformed into spherical coordinates that are divided into multiple spherical regions, and an auxiliary classification task predicts which spherical region each point belongs to, guiding the regression process. Additionally, we incorporate a dynamic weighting scheme and a style pooling layer with multi-head self-attention to capture spectral and temporal dynamics, further boosting performance. This combined training strategy reinforces structured learning and improves prediction consistency. Experimental results show that our approach exceeds baseline methods, confirming the validity of the proposed framework.
@@ 1 Introduction
Speech emotion recognition (SER) predicts a speaker’s emotional state using acoustic features [1]. Understanding the emotional state of speech is important in applications such as human-computer interaction, virtual assistants, mental health monitoring, and conversational artificial intelligence of advances in deep learning [2, 3, 4, 5].
@@ 2.1.1 Normalization and spherical transformation
The original VAD annotations are provided on a $[1,7]$ scale. To facilitate coordinate transformation, we first normalize each of the three emotion dimensions to the range $[-1,1]$. Once normalized, the three-dimensional VAD vector is converted from Cartesian coordinates to spherical coordinates. This transformation decomposes the vector into a magnitude of $r$, azimuth of $\phi$, and elevation of $\theta$.
@@ 2.2.1 Pre-trained SSL model
We adopt WavLM [20] as our feature encoder for extracting features from audio. WavLM is a pre-trained SSL model trained on large-scale data, and it is capable of extracting frame-level high-dimensional representations. To further enhance the model’s ability to capture emotional nuances in speech, we fine-tuned WavLM specifically for emotion recognition.
@@ 2.2.3 Prediction head
From the utterance-level representations generated by style pooling layer, we employ two distinct prediction layers to forecast both the spherical region and the continuous VAD values. To predict the spherical region, a classification layer maps the learned representations to one of the ${N}$ discrete categories. Simultaneously, a regression layer is implemented to output three scalar values, each corresponding to one of the VAD dimensions.
@@ 2.3.1 Spherical region classification loss
To facilitate the alignment of the utterance-level representations with the ${N}$ discrete spherical regions, we empoly an auxiliary classification loss, denoted as $\mathcal{L}_{\text{sph}}$. This loss is computed via a weighted cross-entropy (WCE) loss between the predicted and ground-truth spherical region labels, as the following equation:
@@ 3.1 Experimental setup
We utilized the MSP-Podcast corpus dataset [30], which comprises approximately 237 hours of speech data annotated with both categorical emotion labels and dimensional VAD values. Each utterance has been perceptually annotated by at least five raters. The dataset includes eight categorical emotion classes: happiness, sadness, fear, surprise, contempt, disgust, and a neutral state. For dimensional emotion labels, raters evaluated arousal, dominance, and valence using a seven-point Likert scale. We adhered to the challenge organizers’ guidelines for data partitioning, utilizing the recommended training split for model development [31]. Additionally, to ensure a balanced representation of categories within the validation set, we further divided the original validation set into a validation set and an in-house test set. Specifically, the validation set was constructed by sampling 300 instances per category, maintaining uniform class distribution, while the remaining data was designated as the in-house test set. The “X” label corresponds to instances where no plurality voting winner could be determined among the emotion categories, reflecting cases of annotator disagreement or ambiguity. To minimize label noise and ensure more reliable supervision, all instances with the “X” label were excluded from the training, validation, and in-house test sets.
@@ 3.4 Model performance
As shown in Table 2.2.3, our model improves performance, explained as follows: 1) “w/o Data Preprocessing” indicates that the model is trained on the entire dataset without excluding category data with the “X” label. By removing instances with multiple speakers or ambiguous emotions, the model achieved better performance in predicting emotional attributes. 2) “w/o Style Pooling Layer” denotes to attentive statistics pooling [34], which employs an attention mechanism to assign different weights to frames. By replacing attentive statistics pooling with the style pooling layer, our model benefits from a more expressive style representation. Unlike traditional pooling methods, the style pooling layer effectively captures both spectral and temporal dynamics while leveraging multi-head self-attention to enhance feature extraction. This allows the model to generate richer and more speaker-adaptive embeddings, ultimately leading to improved performance in predicting emotional attributes. 3) “w/o Auxiliary Loss” indicates the absence of an additional auxiliary loss designed to enhance VAD prediction. Specifically, we introduce an auxiliary loss that converts the original VAD values into a spherical emotion vector representation, predicting its corresponding region in spherical coordinates using WCE as a categorical loss. This auxiliary objective aids the model in the early stages of training by providing additional supervision, ultimately leading to improved VAD prediction performance. These results confirm that the proposed method effectively enhances the prediction of emotional attributes and demonstrates the ability of the SER model to adapt to various emotions and speaker styles in the prediction of VAD.
@@ 3.5 Impact of auxiliary loss modification
To evaluate and compare the effectiveness of our proposed auxiliary loss, we conducted three additional experiments: (1) “w/o Dynamic Weighting” refers to training without dynamically adjusting the auxiliary loss weight, instead maintaining a constant weight throughout the entire training process. (2) “w/ Categorical Recognition” replaces the spherical region classification with a categorical emotion prediction task, where the model uses categorical emotion labels as the auxiliary loss instead. (3) “w/ Cross Entropy” replaces WCE with standard cross-entropy in the spherical region classification task.

### v2

@@ Abstract
Speech emotion recognition predicts a speaker's emotional state from speech signals using discrete labels or continuous dimensions such as arousal, valence, and dominance (VAD). We propose EmoSphere-SER, a joint model that integrates spherical VAD region classification to guide VAD regression for improved emotion prediction. In our framework, VAD values are transformed into spherical coordinates that are divided into multiple spherical regions, and an auxiliary classification task predicts which spherical region each point belongs to, guiding the regression process. Additionally, we incorporate a dynamic weighting scheme and a style pooling layer with multi-head self-attention to capture spectral and temporal dynamics, further boosting performance. This combined training strategy reinforces structured learning and improves prediction consistency. Experimental results show that our approach exceeds baseline methods, confirming the validity of the proposed framework.
@@ 1 Introduction
Speech emotion recognition (SER) predicts a speaker's emotional state using acoustic features [1]. Understanding the emotional state of speech is important in applications such as human-computer interaction, virtual assistants, mental health monitoring, and conversational artificial intelligence of advances in deep learning [2, 3, 4, 5].
@@ 2.1.1 Normalization and spherical transformation
The original VAD annotations are provided on a $[1,7]$ scale. To facilitate coordinate transformation, we first normalize each of the three emotion dimensions to the range [1,1][-1,1]. Once normalized, the three-dimensional VAD vector is converted from Cartesian coordinates to spherical coordinates. This transformation decomposes the vector into a magnitude of $r$, azimuth of $\phi$, and elevation of $\theta$.
@@ 2.2.1 Pre-trained SSL model
We adopt WavLM [20] as our feature encoder for extracting features from audio. WavLM is a pre-trained SSL model trained on large-scale data, and it is capable of extracting frame-level high-dimensional representations. To further enhance the model's ability to capture emotional nuances in speech, we fine-tuned WavLM specifically for emotion recognition.
@@ 2.2.3 Prediction head
From the utterance-level representations generated by style pooling layer, we employ two distinct prediction layers to forecast both the spherical region and the continuous VAD values. To predict the spherical region, a classification layer maps the learned representations to one of the N{N} discrete categories. Simultaneously, a regression layer is implemented to output three scalar values, each corresponding to one of the VAD dimensions.
@@ 2.3.1 Spherical region classification loss
To facilitate the alignment of the utterance-level representations with the N{N} discrete spherical regions, we empoly an auxiliary classification loss, denoted as $\mathcal{L}_{\text{sph}}$. This loss is computed via a weighted cross-entropy (WCE) loss between the predicted and ground-truth spherical region labels, as the following equation:
@@ 3.1 Experimental setup
We utilized the MSP-Podcast corpus dataset [30], which comprises approximately 237 hours of speech data annotated with both categorical emotion labels and dimensional VAD values. Each utterance has been perceptually annotated by at least five raters. The dataset includes eight categorical emotion classes: happiness, sadness, fear, surprise, contempt, disgust, and a neutral state. For dimensional emotion labels, raters evaluated arousal, dominance, and valence using a seven-point Likert scale. We adhered to the challenge organizers' guidelines for data partitioning, utilizing the recommended training split for model development [31]. Additionally, to ensure a balanced representation of categories within the validation set, we further divided the original validation set into a validation set and an in-house test set. Specifically, the validation set was constructed by sampling 300 instances per category, maintaining uniform class distribution, while the remaining data was designated as the in-house test set. The ``X'' label corresponds to instances where no plurality voting winner could be determined among the emotion categories, reflecting cases of annotator disagreement or ambiguity. To minimize label noise and ensure more reliable supervision, all instances with the ``X'' label were excluded from the training, validation, and in-house test sets.
@@ 3.4 Model performance
As shown in Table 1, our model improves performance, explained as follows: 1) ``w/o Data Preprocessing'' indicates that the model is trained on the entire dataset without excluding category data with the ``X'' label. By removing instances with multiple speakers or ambiguous emotions, the model achieved better performance in predicting emotional attributes. 2) ``w/o Style Pooling Layer'' denotes to attentive statistics pooling [34], which employs an attention mechanism to assign different weights to frames. By replacing attentive statistics pooling with the style pooling layer, our model benefits from a more expressive style representation. Unlike traditional pooling methods, the style pooling layer effectively captures both spectral and temporal dynamics while leveraging multi-head self-attention to enhance feature extraction. This allows the model to generate richer and more speaker-adaptive embeddings, ultimately leading to improved performance in predicting emotional attributes. 3) ``w/o Auxiliary Loss'' indicates the absence of an additional auxiliary loss designed to enhance VAD prediction. Specifically, we introduce an auxiliary loss that converts the original VAD values into a spherical emotion vector representation, predicting its corresponding region in spherical coordinates using WCE as a categorical loss. This auxiliary objective aids the model in the early stages of training by providing additional supervision, ultimately leading to improved VAD prediction performance. These results confirm that the proposed method effectively enhances the prediction of emotional attributes and demonstrates the ability of the SER model to adapt to various emotions and speaker styles in the prediction of VAD.
@@ 3.5 Impact of auxiliary loss modification
To evaluate and compare the effectiveness of our proposed auxiliary loss, we conducted three additional experiments: (1) ``w/o Dynamic Weighting'' refers to training without dynamically adjusting the auxiliary loss weight, instead maintaining a constant weight throughout the entire training process. (2) ``w/ Categorical Recognition'' replaces the spherical region classification with a categorical emotion prediction task, where the model uses categorical emotion labels as the auxiliary loss instead. (3) ``w/ Cross Entropy'' replaces WCE with standard cross-entropy in the spherical region classification task.